eiffel tower is 497 ft. taller than the washington monument. if each of the monuments were 58 ft. shorter, the eifel tower would be twice as tall as the washingtom monument. how tall is each
Answer by scott8148(6628)

e = w + 497

e - 58 = 2(w - 58)

substitute the equivalent value of e from the 1st equation into the 2nd equation to find w

substitute the numerical value of w into the 1st equation to find e